HUBBLE SPACE TELESCOPE DAILY REPORT #1370 PERIOD COVERED: 0000Z (UTC) 04/14/95 - 0000Z (UTC) 04/17/95 Daily Status Report as of 107/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1 5162 (A High Resolution UV Spectral Survey of the Sharp Line Star Chi-Lupi The High Resolution Spectrograph (Side One) was used to make ultraviolet Echelle spectrographic observations of the sharp lined chemically peculiar star HD141556 (aka, Chi Lupi). The objective is to obtain a spectroscopic data base for use in element abundance analysis, and as a standard of reference for the study of normal and peculiar B and A class stars. GHRS observations #65 was reported to have a carrousel reset failure at 104/0440Z. A routine status buffer error message was received (HSTAR 5180). No other anomalies were noted. 1.2 Completed FGS 6051 (Dynamical Masses For The Stars in the Pre-Main Sequence Spectroscopic Class) The FGS were used to make astrometric measurements of the location of star STAR-04562 (aka, GSC2387+535) to measure its relative orbit in a pre-main sequence binary system. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 5672 (Ages For The Outermost Globular Clusters) The WF/PC-2 was used to make both deep and short exposure observations of the central regions of globular cluster PAL3-Centaurus. This cluster was chosen as it represents an early staging in the formation of the galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.5 Completed Five Sets of WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of TOL1059+10, NGC-3862, NGC-4507, HEAO0510+0 and NGC-7552 to study the interactions between the central engine, the stars, and the interstellar matter in these AGNs. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of FOS/BL 5723 (Constraints of Massive Star Evolution From Carbon and Oxygen Abundance) The Faint Object Spectrograph (Blue Side) made ultraviolet spectrographic observations of the massive Wolf-Rayet star HD32402 to study its Carbon and oxygen abundances. The observations were completed as planned and no anomalies were reported. 1.7 Completed GHRS/1 5412 (The Boundary Layer White Dwarf Connection: CV White Dwarf Rotation) The High Resolution Spectrograph (Side One) was used to make observations of the white dwarf nova star U-Gem after a period of outbursts. The observations were completed as scheduled, and no anomalies were reported. 1.8 Completed WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxy NGC-4414 to be used in measuring the periods of Cepheid variable stars within this galaxy. The observations were completed as planned, and no anomalies were noted. 1.9 Completed Two Sets of WF/PC-2 5167 (Jovian Auroral Lyman-Alpha Profile: Cycle 4 ) The WF/PC-2 was used to make observations of the northern auroral regions of the planet Jupiter to study the excitation processes and atmospheric properties in this region. The observations were completed as planned, and no problems were reported. 1.10 Completed FOS/BL 6146 (FOS Location of Spectra and Aperture - Part 2) The Faint Object Spectrograph (Blue Side) was used to conduct a test to determine the repeatability of positioning the filter-grating wheel. All observations were internal. Wavelength and Dark calibration images were also taken. All observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 5512 (Cores of Early Type Galaxies: Part II) The WF/PC-2 was used to make observations of the nucleus of early type galaxy NGC-3377-NU. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6142 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Smaller Set Part II) The WF/PC-2 was used to make Earth Flat calibration images for the ultraviolet and narrow-band filters to correct for the telescope's illumination pattern. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed FOC/96 5153 (UV-Bright Stars in Dwarf Galaxy Leo I) The Faint Object Spectrograph (f/96) was used to make ultraviolet observations of UV bright start in the dwarf galaxy Leo I using several different bypass filters. The intent is to search for a stellar UV flux source that may account for the UV flux observed in normal galaxies. The observations were executed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 18 Successful : 18 Scheduled Reacquisitions : 24 Successful : 24 2.2 FHST Updates: Scheduled : 28 Successful : 28 2.3 Significant HSTARs: None 2.4 Operations Notes: There was only one data inversion this past week-end at STGT. The problem appears to be fixed with new firmware. Engineers are looking closely at the FOS for possible microprocessor reset problems. At 0517Z today, while the GHRS was outputting science data at 4k, the FOS went from low voltage to high voltage on the Blue Side. This voltage transition evidently caused a spontaneous FOS microprocessor reset. These resets have been seen several times before associated with exactly this transition. Immediately after the time of the high voltage transition, a series of status buffer messages were received. The final effect of these events was to disable the GHRS SD interface. It is not yet known to what extent the FOS and GHRS observations at the indicated time were affected. The next FOS Blue Side observation does not occur for some time, however, the GHRS has a condition of the same proposal roughly 1556Z today. On Saturday during a type-2 slew, bad counts were noticed for the HGA spline. The counts quickly returned to normal, and no impact to on-going operations was reported. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-